log(x)^2=2logx This one is also true right? Thank you for you help.
Found 2 solutions by stanbon, Nate:
Answer by stanbon(48535)

Yes.
Cheers,
Stan H.

Answer by Nate(3500)

the power property of logarithms shows:
log (x^a) = a log x
so....
log (x)^2 = 2 log (x) is the same as:
log (x^2) = 2 log (x) it is true